# On $\mathcal{F}$-semi-transitive families of operators

### v6

@@ 5.1. Main results
5.2. Applications
We will now apply our results to the case of generalized weighted bilateral shift on $\ell_{2}(\mathcal{C}).$ More precisely, put $\mathcal{A}=\ell_{2}(\mathcal{C}),$ and
@@ 5.2. Applications
For $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$ and let $\mathcal{U}_{1},\dots,\mathcal{U}_{N}$ are unitary operators on $H$. For each $n\in\mathbb{N}$ and $s\in\{1,\ldots,N\},$ we define the map $\Phi_{n,s}:\mathcal{A}_{1}\to\mathcal{A}_{1}$ by
$\Phi_{n,s}\left(\{y_{j}\}_{j\in\mathbb{Z}}\right)=\{\mathcal{U}_{s}^{-r_{s}n}y_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}},$
for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1},$ (that is the $j$-th coordinate of $\Phi_{n,s}(\{y_{j}\}_{j\in\mathbb{Z}})=\mathcal{U}_{s}^{-r_{s}n}y_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}$ for all $j\in\mathbb{Z}$).
Now, the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\mathcal{A}=\ell_{2}(\mathcal{C})$ and $\tilde{p}_{J,m}=\tilde{p}_{J,m}^{3}$ for all $(J,m)\in\mathbb{N}^{2}.$ If $\mathcal{F}$ is a family of subsets of $\mathbb{N}$ which is finitely invariant, that is, for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$ it holds that $F\setminus\{1,\dots,n\}\in\mathcal{F}$ also, then, by some calculations it can be checked that $\{\Phi_{n,1}\}_{n\in\mathbb{N}},\cdots,\{\Phi_{n,N}\}_{t\in\mathbb{N}}$ is a disjoint aperiodic system with respect to the left approximate unit $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ and that the condition (R) is satisfied in this case.
For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ let $\{W_{n,j}^{(\ell)}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm and such that each $W_{n,j}^{(\ell)}$ has a bounded inverse. Then, for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, we let $b_{n,\ell}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell})_{i}=W_{n,i}^{(\ell)}\mathcal{U}_{\ell}^{r_{\ell}n},$
for all $i\in\mathbb{Z}.$ Since for each $n\in\mathbb{N},i\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\},$ the operator $W_{n,i}^{(\ell)}\mathcal{U}_{\ell}^{r_{\ell}n}$ is invertible, it follows that the condition (C) is satisfied in this case.
Then we let $T_{n,\ell}$ be the operator on $\mathcal{A}$ given by
$T_{n,\ell}(a)=b_{n,\ell}\,\Phi_{n,\ell}(a),\qquad\text{for all }a\in\mathcal{A}.$
Once again, since $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}=\{\tilde{p}_{J,m}^{3}\}_{(J,m)\in\mathbb{N}^{2}}$ is a left approximate unit for $\mathcal{A}$ (so that this left approximate unit is indexed over $\mathbb{N}^{2}$), by some calculations one can check that for all $n,J,m\in\mathbb{N},\,s\in\{1,\dots,N\}$ it holds that
$(b_{n,s,(J,m)}^{-1})_{i}=(b_{n,s,J}^{-1})_{i}=\mathcal{U}_{s}^{-r_{s}n}\,W^{(s)\,-1}_{n,i},$
for all $i\in\mathbb{Z}$ with $J\leq i\leq J,$ and $(b_{n,s,(J,m)}^{-1})_{i}=(b_{n,s,J}^{-1})_{i}=0$ for all $i\in\mathbb{Z}$ with $i\leq-J-1$ and all $i\in\mathbb{Z}$ with $i\geq J+1.$ In fact, we write just $b_{n,s,J}^{-1}$ instead of $b_{n,s,(J,m)}^{-1}$ because it does only depend on $J$ and not on $m.$
Finally, note also that for each $a\in\mathcal{A}$ and $j\in\mathbb{Z}$ it holds that $\|a_{j}\|\leq\|a\|.$
Taking all this into account, it is not hard to deduce Theorem 4.2 as a special case of Theorem 5.1.
From now on and in the rest of this section we will assume that $\mathcal{F}$ the family of all infinite subsets of $\mathbb{N}.$ In view or Corollary 5.2, by arguments similar to those in the proofs of [27, Theorem 3.2] and [28, Proposition 2.7], it is neither hard to deduce the following proposition in this special case.
@@
Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
$\lim_{k\to\infty}\big\|W_{n_{k},j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\big\|W_{n_{k},j+r_{\ell}n_{k}}^{(\ell)}\,x^{(j)}\big\|=0,$
and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
$\lim_{k\to\infty}\|W_{n_{k},j-r_{s}n_{k}+r_{\ell}n_{k}}^{(\ell)}\,W_{n_{k},j}^{(s)-1}y_{s}^{(j)}\|=0,$
then the sequences of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Next we let $\Omega$ be a locally compact Hausdorff space, $\alpha_{1},\dots,\alpha_{N}$ be homeomorphisms of $\Omega$, $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\Omega)$, where $C_{b}(\Omega)$ denotes the space of all bounded continuous functions on $\Omega$. We will assume that $w_{j}^{(\ell)}>0$ for all $j\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$ and that $\|w_{j}^{(\ell)}\|_{\infty}<M$ for all $j\in\mathbb{Z}$, $\ell\in\{1,\dots,N\}$ and some $M>0.$ Now we consider $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$, and
@@
For each pair of compact subset $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$. For $J\in\mathbb{N}$, we let $\tilde{p}_{(K_{1},K_{2},J)}\in\mathcal{A}$ be given by
@@
$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha_{s}^{r_{s}n_{k}-i}\big)(t)}\Bigg)\Bigg(\sup_{t\in K}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha_{\ell}^{-i}\big)(t)\Bigg)=0$

### v7

@@ 2. Main results
Now we set $N=1$ and we will therefore write $b_{t}$, $\Phi_{t}$ and $T_{t}$ instead of $b_{t,1}$, $\Phi_{t,1}$ and $T_{t,1},$ respectively. We will now assume that $b_{t}$ is invertible for each $t\in\mathcal{S}$ and we will consider the map $S_{t}:\mathcal{A}\rightarrow\mathcal{A}$ given by $S_{t}(a)=\Phi_{t}^{-1}(b_{t}^{-1})\,\Phi_{t}^{-1}(a)$ for all $a\in\mathcal{A}$. Obviously, $S_{t}=T_{t}^{-1}$ for each $t\in\mathcal{S}.$
We set $C_{t}:=\frac{1}{2}(T_{t}+S_{t})$ for all $t\in\mathcal{S}$.
Proposition 2.4.
(1) The family $\{C_{t}\}_{t\in\mathcal{S}}$ is $\mathcal{F}$–semi–transitive.
(2) For each fixed $p_{\alpha}$ and $\varepsilon>0$ there exists some $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t}\}_{t\in F}$, $\{e_{t}\}_{t\in F}$, $\{f_{t}\}_{t\in F}$ with $e_{t}+f_{t}=g_{t}$ for all $t\in F$ such that
$\|d_{t}-p_{\alpha}^{2}\|<\varepsilon,\qquad\|g_{t}-p_{\alpha}^{2}\|<\varepsilon,$
$\|\Phi_{t}^{-2}(b_{t})\Phi_{t}^{-1}(b_{t})e_{t}\|<\varepsilon,$
$\|\Phi_{t}(b_{t}^{-1})\,b_{t}^{-1}\,f_{t}\|<\varepsilon,$
$\|\Phi_{t}^{-1}(b_{t})\,d_{t}\|\,\|\Phi_{t}^{-1}(b_{t})\,e_{t}\|<\varepsilon^{2},$
$\|\Phi_{t}^{-1}(b_{t})\,d_{t}\|\,\|b_{t}^{-1}f_{t}\|<\varepsilon^{2},$
$\|b_{t}^{-1}\,d_{t}\|\,\|\Phi_{t}^{-1}(b_{t})\,e_{t}\|<\varepsilon^{2},$
$\|b_{t}^{-1}\,d_{t}\|\,\|b_{t}^{-1}\,f_{t}\|<\varepsilon^{2},\text{ for all }t\in F.$
Proof.
Given non–empty open subsets $\mathcal{O}$ and $\mathcal{V}$ of $\mathcal{A}$, as in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$ we can find some $\delta>0$ and some $p_{\alpha}^{3}x\in\mathcal{O}\setminus\{0\}$, $p_{\alpha}^{3}y\in\mathcal{V}\setminus\{0\}$ such that $\delta$–neighbourhoods of $p_{\alpha}^{3}x$ and $p_{\alpha}^{3}y$ are contained in $\mathcal{O}\setminus\{0\}$ and $V\setminus\{0\}$, respectively.
Set
$\varepsilon:=\min\left\{\frac{\delta}{4\|p_{\alpha}\|(\|x\|+\|y\|)},\;\frac{\delta}{16\|p_{\alpha}\|\sqrt{(\|x\|\|y\|)}},\;\frac{\|p_{\alpha}^{3}x\|}{2\|p_{\alpha}x\|},\;\frac{\|p_{\alpha}^{3}y\|}{2\|p_{\alpha}y\|}\right\}.$
Choose $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t}\}_{t\in F}$ , $\{e_{t}\}_{t\in F}$, $\{f_{t}\}_{t\in F}$ satisfying the assumptions in (2) with respect to $p_{\alpha}$ and $\varepsilon$. Since $e_{t}+f_{t}=g_{t}$ and $g_{t}\,p_{\alpha}y\neq 0$ (as shown in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$), it follows that either $e_{t}\,p_{\alpha}y\neq 0$ or $f_{t}\,p_{\alpha}y\neq 0$. Then, by the same arguments as in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$ we can show that either $T_{t}(e_{t},p_{\alpha}y)\neq 0$ or $S_{t}(f_{t},p_{\alpha}y)\neq 0$.
Further, we notice that for each $t\in F$ it holds that
$\|T_{t}^{2}(e_{t},p_{\alpha}y)\|=\|b_{t}\,\Phi_{t}\,(b_{t})\,\Phi_{t}^{2}\,(e_{t}\,p_{\alpha}y)\|$
$\leq\|\Phi_{t}^{-2}\,(b_{t})\,\Phi_{t}^{-1}\,(b_{t})\,e_{t}\|\,\|p_{\alpha}\|\,\|y\|$
and,
$\|S_{t}^{2}(f_{t},p_{\alpha}y)\|=\|\Phi_{t}^{-1}(b_{t}^{-1})\cdot\Phi_{t}^{-2}(b_{t}^{-1}f_{t}\,p_{\alpha}y)\|\leq\|\Phi_{t}(b_{t}^{-1})b_{t}^{-1}f_{t}\|\,\|p_{\alpha}\|\,\|\|y\|.$
For each $t\in F,$ put
$v_{t}:=d_{t}p_{\alpha}x+\frac{2\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}{\sqrt{\|T_{t}(e_{t}+p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}(T_{t}(e_{t}+p_{\alpha}x)+S_{t}(e_{t}p_{\alpha}x)).$
By the above inequalities and inequalities in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$, the choice of $\varepsilon$ and the triangle inequality, we can deduce that $v_{t}\in\mathcal{O}$ and
$\frac{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}{\sqrt{\|T_{t}(d_{t}+p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}C_{t}(v_{t})\in\mathcal{V}\text{ for all }t\in F.$
By putting either $e_{t}=0$ (so that $f_{t}=g_{t}$), or by putting $f_{t}=0$ (so that $e_{t}=g_{t}$) in Proposition 2.4, then we can get the following corollary.
Corollary 2.5.
Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}$. If there exist dense subsets $\mathcal{B}_{1},\mathcal{B}_{2}$ of $\mathcal{A}$ and a strictly increasing sequence $\{n_{k}\}\subseteq\mathbb{N}$ such that for every fixed $p_{\alpha}$ it holds that either and
$\lim_{k\to\infty}\|\Phi_{n_{k}}^{-2}(b_{n_{k}})\,\Phi_{n_{k}}^{-1}(b_{n_{k}})\,y\|=\lim_{k\to\infty}\|b^{-1}_{n_{k}}x\|\,\|\Phi_{n_{k}}^{-1}(b_{n_{k}})y\|=$
$\lim_{k\to\infty}\|\Phi_{n_{k}}^{-1}(b_{n_{k}})x\|\,\|\Phi_{n_{k}}^{-1}(b_{n_{k}})y\|=0$
or
$\lim_{k\to\infty}\|\Phi_{n_{k}}(b_{n_{k}}^{-1})b_{n_{k}}^{-1}y\|=\lim_{k\to\infty}\|b_{n_{k}}^{-1}x)\|\cdot\|b_{n_{k}}^{-1}y\|=\lim_{k\to\infty}\|\Phi_{n_{k}}^{-1}(b_{n_{k}})\|\|b_{n_{k}}^{-1}y\|=0$
for all $x\in\mathcal{B}_{1}$ and $y\in\mathcal{B}_{2}$, then the family $\{C_{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$–semi-transitive.
3. Applications
From now on and in the rest of this section we will assume that $\mathcal{F}$ the family of all infinite subsets of $\mathbb{N}.$ Further, in this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$. For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$.
Moreover, we will denote by $\Omega$ be a locally compact Hausdorff, non-compact space. As usual, $C_{b}(\Omega)$ will denote the space of all bounded continuous functions on $\Omega$ and $C_{0}(\Omega)$ will denote the space of all continuous functions on $\Omega$ vanishing at infinity. Both these spaces will be equipped with supremum norm. In addition, we let $C_{c}(\Omega)$ is the space of all continuous compactly supported functions on $\Omega.$ For each pair of compact subset $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$.
3.1. Applications to generalized weighted shifts on Hilbert C*-modules
In this subsection, we will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [21, Example 1.3.5]. Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra. Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
$\sum_{j}y_{j}^{*}x_{j}^{*}x_{j}y_{j}\leq\sum_{j}y_{j}^{*}\|x_{j}\|^{2}y_{j}\leq\parallel\{x_{j}\}_{j\in\mathbb{Z}}\parallel^{2}\sum y_{j}^{*}y_{j}.$
[…]
As shown in the proof of [16, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$. We put then $\mathcal{A}=\ell_{2}(\mathcal{C}),$ and
@@ 3.1. Applications to generalized weighted shifts on Hilbert C*-modules
Let now $\mathcal{U}$ be a unitary operator on $H$. For each $n\in\mathbb{N}$ we define the map $\Phi_{n}:\mathcal{A}_{1}\to\mathcal{A}_{1}$ by
$\Phi_{n}\left(\{y_{j}\}_{j\in\mathbb{Z}}\right)=\{\mathcal{U}^{-n}y_{j-n}\mathcal{U}^{n}\}_{j\in\mathbb{Z}},$
for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1},$ (that is the $j$-th coordinate of $\Phi_{n}(\{y_{j}\}_{j\in\mathbb{Z}})=\mathcal{U}^{-n}y_{j-n}\mathcal{U}^{n}$ for all $j\in\mathbb{Z}$).
Now, the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\mathcal{A}=\ell_{2}(\mathcal{C})$ and $\tilde{p}_{J,m}=\tilde{p}_{J,m}^{3}$ for all $(J,m)\in\mathbb{N}^{2}.$ Moreover, by some calculations it can be checked that $\{\Phi_{n}\}_{n\in\mathbb{N}}$ is aperiodic with respect to the left approximate unit $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ and that the condition (R) is satisfied in this case.
For each $n\in\mathbb{N},$ let $\{W_{n,j}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm such that each $W_{n,j}$ has a bounded inverse and such that $\{W_{n,j}^{-1}\}_{j\in\mathbb{Z}}$ is also uniformly bounded in the norm. Then, for each $n\in\mathbb{N}$ we let $b_{n}\in\mathcal{A}_{1}$ be given by
$(b_{n})_{i}=W_{n,i}\mathcal{U}^{n},$
for all $i\in\mathbb{Z}.$ Obviously, $b_{n}$ is invertible in $\mathcal{A}_{1}$ for each $n\in\mathbb{N}$ and
$(b_{n}^{-1})_{i}=\mathcal{U}^{-n}\,W^{\,-1}_{n,i},$
for all $i\in\mathbb{Z},$ hence, the condition $(C)$ is satisfied in this case.
Finally, for each $n\in\mathbb{N}$ we let $T_{n}$ be the operator on $\mathcal{A}$ given by
$T_{n}(a)=b_{n}\,\Phi_{n}(a),\qquad\text{for all }a\in\mathcal{A}.$
Noticing that for each $a\in\mathcal{A}$ and $j\in\mathbb{Z}$ it holds that $\|a_{j}\|\leq\|a\|$, in view of Corollary 2.5 and by arguments similar to those in the proofs of [16, Theorem 3.2] and [17, Proposition 2.7] it is not hard to deduce the following proposition in this special case.
@@
Under the above notation and assumptions, if there exist dense subsets $H_{0}$ and $H_{1}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}$ of natural numbers such that either
$\lim_{k\to\infty}\|W^{-1}_{n_{k},j}x\|\,\|W_{n_{k,j+n_{k}}}y\|=\lim_{k\to\infty}\|W_{n_{k},j+n_{k}}x\|\,\|W_{n_{k},j+n_{k}}y\|$
$\lim_{k\to\infty}\|W_{n_{k},j+2n_{k}}W_{n_{k},j+n_{k}}y\|=0,$
or
$\lim_{k\to\infty}\|W_{n_{k},j+n_{k}}x\|\,\|W^{-1}_{n_{k},j}y\|=\lim_{k\to\infty}\|W^{-1}_{n_{k},j}x\|\,\|W^{-1}_{n_{k},j}y\|=$
$\lim_{k\to\infty}\|W^{-1}_{n_{k},j-n_{k}}W^{-1}_{n_{k},j}y\|=0,$
for all $x\in H_{0}$ and $y\in H_{1}$, then $\{C_{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$–semi-transitive on $\mathcal{A}$.
Example 3.2.
Let $H=L^{2}(\mathbb{R})$. As in [16, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel,\parallel w_{j}^{-1}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq K$ for all $j\in\mathbb{Z}$ and some $K>0$. For each $j\in\mathbb{Z},$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R},$ and let then $V_{j}\in B(H)$ be given by
$V_{j}(f)=w_{j}\cdot(f\circ\alpha_{j})$
for all $f\in H.$ For each $n\in\mathbb{N},i\in\mathbb{Z}$, let $W_{n,i}\in B(H)$ be given by
$W_{n,i}:=V_{i}V_{i-1}\cdots V_{i-n+1}.$
If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [16, Example 3.3] one can show that the conditions of Proposition 3.1 are satisfied with $H_{0}=H_{1}=C_{c}(\mathbb{R})$ in this case. On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 3.1 will be satisfied with $H_{0}=H_{1}=C_{c}(\mathbb{R}).$
Next, for $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}.$
Let $\alpha_{1},\dots,\alpha_{N}$ be homeomorphisms of $\Omega$, $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\Omega)$. We will assume that $w_{j}^{(\ell)}>0$ for all $j\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$ and that $\|w_{j}^{(\ell)}\|_{\infty}<M$ for all $j\in\mathbb{Z}$, $\ell\in\{1,\dots,N\}$ and some $M>0.$ Now we consider $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$, and
@@
For $J\in\mathbb{N}$, we let $\tilde{p}_{(K_{1},K_{2},J)}\in\mathcal{A}$ be given by
@@
$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha_{s}^{r_{s}n_{k}-i}\big)(t)}\Bigg)\cdot\Bigg(\sup_{t\in K}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha_{\ell}^{-i}\big)(t)\Bigg)=0$